# New Coevolution Dynamic as an Optimization Strategy in Group Problem Solving

### v2

@@ 3.3 Dynamics with coevolution
To enhance the effect of coevolution, we restricted the values of the average degree for the analysis. We observed different scenarios for different values of this degree: coevolution works for some cases and does not work at other cases. We say it works when the computational cost $\left\langle C\right\rangle$ is lower compared to the Reference Case, keeping all other parameters the same. We found local improvements on $\left\langle C\right\rangle$ at the degrees $\mu=1.0$ and $\mu=10.0$ (see Fig. 3). These are the scenarios where coevolution works (at some values of $M$), depicted in Figs. 4, 5 and 8(a). The scenarios where it does not work are depicted in Figs. 8(b) and 8(c).
Let’s start with the low average degree scenario, Fig. 3(a). Considering the easy landscape with $K=0$ (Fig. 4(a)), the coevolution brings benefit to the cost except when $L=1$. As there is only one maximum, any other agent with a higher fitness is worth copying. And this process is facilitated by the selective rewiring of the coevolution. Interestingly, the case $L=1$ increases the cost. The difficult case $K=4$ (Fig. 4(b)) has some similarities: the cost increases at $L=1$ and then decreases up to $L=5$. At $L=6$ there is a significant increase only at $M\sim 60$.
At higher average degree (Fig. 3(b)) the results are drastically different. The ease case $K=0$ (Fig. 5(a)) shows influence of $L$ only on small network sizes around $M\lesssim 15$. At $K=4$ the results become more interesting. At small networks, coevolution continues to present a higher computational cost. However, at medium-sized networks ($20\leq M\leq 120$), we see a significant improvement in computational cost: $\left\langle C\right\rangle$ presents a maximum at $M=67$ for $L=0$ (Fig. 5(b)) and a minimum at $L=3$.
Figure 5: Coevolution results in a high degree network ($\mu=10$) on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of modification $L$, limited by $L<k_{i}$ for each agent. $L=0$ means no coevolution. Parameters: $S=10^{5}$ and $\mu=10$. The green line is the IS (Eq. 4) and the black dashed line is the FCN result. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.
The minimum of $\left\langle C\right\rangle$ with respect to $L$ has a larger sensitivity to $M$ than with to $\mu$. As can be seen on Fig. 6(a), the cost has a pronounced minimum between $L=2$ and 3 considering different values of $M$. However, when $M$ is fixed and $\mu$ is varied, the minimum is not so evident (Fig. 6(b)), and even not present at some values.
Figure 6: Characterization of the computational cost minimum caused by the coevolution. Computational cost $\left\langle C\right\rangle$ vs. $L$. Parameters: $K=4$, $q=0$ and $S=30000$. (a) Minimum on the $(M,L)$ space. Different values of $M$ while $\mu=10$. (b) Different values of $\mu$ while $M=67$.
@@
Figure 7: 2D map of the computational cost $\left\langle C\right\rangle$ and its gain $\Delta$ in the $(L,\mu)$ space. Parameters: $K=4$, $q=1$, $M=67$ and $S=10^{5}$. (a) $\log\left\langle C\right\rangle$. (b) $\Delta$ defined on Eq. 8. Positive values are indicated in red and negative values in blue.
3.4 Other dynamics
We now verify the differences between the coevolution ($q=1$) and constant rewiring ($q=0$) dynamics. Considering the curve $L=1$ from Fig. 5(b), the constant rewiring case gives a lower cost, as can be seen on Fig. 8(a). Focusing on the peak at $M\sim 67$, the results shows a slightly better performance (lower cost) with $L=1$ for the constant rewiring case, while negligible differences when $L=3$. Figures 8(b) and 8(c) show these same results for different average degrees $\mu$, where coevolution itself does not work. One can see on these results that constant rewiring does not work either.
@@ 3.4 Other dynamics
Figure 8: Comparison between coevolution and constant rewiring dynamics. Computational cost $\left\langle C\right\rangle$ as a function of network size $M$ for different values of the average degree. The green line is the IS (Eq. 4), the black dashed line is the FCN result and RC is the reference case (equivalent to $L=0$). (a) Scenarion where coevolution works: average degree $\mu=10$. Scenarios where coevolution does not work: (b) $\mu=5$ and (c) $\mu=15$.
Lastly we studied the impact of non-linearity on the computational cost, which happens when $q$ is different from 0 or 1 [38]. In this work non-linearity yield no significant outcome, as can be shown on Fig. 9. The case $q=0.5$ only showed slightly better performance than the linear case $q=1$ around $M=67$. Other than this region there were no significant differences.
Figure 9: Study of the impact of non-linear effects on the coevolution dynamics. Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of the non-linear parameter $q$. Parameters: $\mu=10$, $S=3000$ and $L=3$. The violet circles correspond to $q=0.5$, the orange circles to $q=1.5$, and the red circles to $q=2$. The reference case RC ($L=0$ and $q=1$) is depicted in gray. The non-linear case refers to $q\neq 0.0$ and $q\neq 1.0$.
@@ 4 Conclusions
A natural strategy for the target is to change some of his neighbors if they have worst (smaller) fitness than him. Nonetheless, this rule did not give satisfactory results in our tests when compared with the reference case. Surprisingly, the opposite rule did yield some interesting results. We chose to apply this new coevolution model on a cooperative process. Using as prototype the NK model, our results showed a convoluted behavior of the system efficiency.
At the easy landscape, coevolution did bring significantly better results only at lower average degree (Fig. 4(a)). Additionally, the larger $L$ the lower the $C$, except at $L=1$ at which the results are worse. However, the FCN case is still the more efficient one at $K=0$, regardless other parameters. At the difficult landscape $K=4$, the behaviour was unanticipated. At lower average degree, after a cost increase at $L=1$, the cost cutback for increasing $L$ for most part of the system size range. Except at around $M\sim 70$ where a peak starts to grow at $L=5$ and becomes higher than the IS cost ((Fig. 4(b)). At higher average degree, this peak on the cost already exists at $L=0$ (Fig. 5(b)). It disappears at $L=3$ and comes back at $L=5$. Indeed, this minimum can be better visualized on the $(M,L)$ space (Fig. 6(a)). Fixing $M=67$, the minimum of the computational cost can be seen in details on the 2D map of Fig. 7(b) (blue region).
This work showed that this tunable coevolution can drastically change the behavior of computational cost. It opens some interesting perspectives such as the impact of this tunable coevolution on the phase diagram involving absorbing phase transition of models such as Axelrod’s, voter model and others.
@@ Acknowledgements
This work received financial support from the brazillian agencies CNPq (project 405508/2021-2) and FAPEG (project no. 401425/2023-1). This research was also supported by LaMCAD/UFG.
@@
The model was implemented using Fortran and the analysis and graphics were developed using Python (packages: Numpy [50], Matplotlib [51], NetworkX [52] and Pandas [53]) and Gnuplot.
@@ A.1 NK landscape
Following the formal definition from Eq. 1, we have $f:\mathscr{D}\rightarrow\mathscr{R}$, where $\mathscr{D}$ is the set of $2^{N}$ binary vectors $\mathcal{X}$ and $\mathscr{R}$ is the real numbers set. The argument of the contributing functions $\Phi_{i}$ is given by Eq. 2. To better understand this, suppose the vector $\mathcal{Y}=(y_{0},y_{1},y_{2},y_{3})=(0,1,1,0)$ on the landscape $N=4$ and $K=2$. There will be $K+1=3$ bits on the argument of each $\Phi_{i}$. The first bit from the i-th contribution is the i-th bit of $\mathcal{Y}$, while the other bits are the $K=2$ neighbors on the right. So, the first contribution of $\mathcal{Y}$ is $\Phi_{0}(y_{0},y_{1},y_{2})=(0,1,1)$. The second contribution will be $\Phi_{1}(y_{1},y_{2},y_{3})=\Phi_{1}(1,1,0)$. On the third contribution $\Phi_{2}$, the third bit will be the fifth one of $\mathcal{Y}$. However, and the string $\mathcal{Y}$ has only $N=4$ bits, so the fifth bit will be the first one. This equivalent to subtract $N$ from the position of the bit: 5-4 = 1. Thus: $\Phi_{2}(y_{2},y_{3},y_{0})=\Phi_{2}(1,0,0)$. In the same way: $\Phi_{3}(y_{3},y_{0},y_{1})=\Phi_{2}(0,0,1)$ Putting together the four contributions, we have:
@@ A.2 Base 2 and 10
Conceptually, the description from previous section is enough. However, computationally, it is compelling to represent the vectors $\mathcal{X}$ and the arguments of the contributions $\Phi_{i}$ as integers. Both of them are binary sequences, 0s and 1s, numbers in base 2, and thus, can be converted to decimal numbers (base 10). In general, considering $N=4$: $(x_{0},x_{1},x_{2},x_{3})=x_{3}\cdot 2^{0}+x_{2}\cdot 2^{1}+x_{1}\cdot 2^{2}+x_{0}\cdot 2^{3}$, starting the count from the right. Thus,

### v3

@@ 4.1 Dynamics with coevolution
To enhance the effect of coevolution, we restricted the values of the average degree network $\mu$ for the analysis. We observed different scenarios for different values of $\mu$: coevolution works for some cases and does not work at other cases. We say it works when the computational cost $\left\langle C\right\rangle$ is lower compared to the Reference Case, independent search (IS) and FCN case, keeping all other parameters the same. We will explore some local improvements on $\left\langle C\right\rangle$, at some values of $M$, at degree $\mu=10.0$.
We begin with the results using $r_{1}$ (Eq. 5) as the rewiring probability, and Fig. 4(a) shows the results for the ease case $K=0$. The results shows influence of $L$ only on small network sizes, around $M\lesssim 15$: the gain on $\left\langle C\right\rangle$ improves with $L$. However, at $M\geq 15$, $L$ has no influence at all. FCN result is still the best one for all $M$. At $K=4$ (Fig. 4(b)) the results become more interesting. At small network size $M$, FCN is still the best result. However, at medium-sized networks ($20\leq M\leq 120$), we see a significant improvement in computational cost with $L=2$ and $3$ around $M\sim 70$: $\left\langle C\right\rangle$ is slightly lower than the IS case. More interestingly, $\left\langle C\right\rangle$ rises back at $L=6$, greater than the $L=0$ value.
Moving on with the rewiring probability $r_{2}$ (Eq. 6), the results for the ease case $K=0$ (Fig. 5(a)) are similar as the the ones with $r_{1}$. On the other hand, results for the difficult case $K=4$, depicted on Fig. 5(b), are even more interesting. As before, there is an improvement on $\left\langle C\right\rangle$, but now the minimum is even lower: $\left\langle C\right\rangle\sim 0.9$. At higher $L$, $\left\langle C\right\rangle$ rises back, although lower than the $L=0$ value.
@@
Figure 7: Coevolution results using $r=1.0$ as rewiring probability (constant rewiring). Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of modification $L$, limited by $L<k_{i}$ for each agent. $L=0$ means no coevolution. Parameters: $S=10^{5}$, $\mu=10$ and $K=4$. The green line is the IS (Eq. 4) and the black dashed line is the FCN result.
4.2 Rewiring probabilities
It is clear from Figs. 4(b) and 5(b) that the best gain on $\left\langle C\right\rangle$ is achieved around $M\sim 67$ with $r=r_{2}$. It indicates that rewiring when $\phi_{t}$ is low is the key mechanism. To better understand this result, we studied other rewiring probabilities. Probability $r_{1}$ is maximum when the target is the model and $r_{2}$ is maximum when the target has the worst fitness on the neighborhood. The natural next step is the constant rewiring case $r=1.0$. Figure 7 shows a similar result as the one using probability $r_{2}$ (compare with Fig. 5(b)). Indeed, $r=1.0$ implies a constant rewiring when $\phi_{t}$ is low, similar to $r=r_{2}$, but it also includes constant rewiring when $\phi_{t}$ is higher. So, define two more rewiring probabilities:
$r_{3}=\left\{\begin{array}[]{rl}0.0,&\text{ if }\phi_{t}<\phi_{a},\\ 1.0,&\text{ if }\phi_{t}\geq\phi_{a},\end{array}\right.$ (9)
$r_{4}=\left\{\begin{array}[]{rl}1.0,&\text{ if }\phi_{t}<\phi_{a},\\ 0.0,&\text{ if }\phi_{t}\geq\phi_{a},\end{array}\right.$ (10)
where $\phi_{a}=(\phi_{s}+\phi_{m})/2$ is the average fitness between the limits of the neighborhood. The probabilities $r_{1}$ to $r_{4}$ are displayed on Fig. 8(a). Probability $r_{3}$ implies rewiring only when $\phi_{t}$ is closer to $\phi_{m}$ (larger than the average $\phi_{a}$). It is an extrapolation of probability $r_{1}$. On the other hand, probability $r_{4}$ implies rewiring only when $\phi_{t}$ is closer to the minimum $\phi_{s}$, an extrapolation of $r_{2}$.
@@ 4.2 Rewiring probabilities
Figure 8: Comparison of the results between different rewiring probabilities. (a) Behaviour of four different definitions: $r_{1}$ (Eq. 5), $r_{2}$ (Eq. 6), $r_{3}$ (Eq. 9) and $r_{4}$ (Eq. 10). (b) Computational cost $\left\langle C\right\rangle$ vs. $M$ for 6 possible rewiring probabilities at $L=3$. The black dashed line is the FCN result and the vertical dotted green line indicates the position $M=67$. $r=0.0$ is the Reference case (no coevolution) and $r=1.0$ is the constant rewiring one. Parameters: $K=4$, $S=10^{5}$ and $\mu=10$. The green line is the IS (Eq. 4).
Figure 9: Comparison of the results between different rewiring probabilities: Computational cost $\left\langle C\right\rangle$ vs. $L$. Parameters: $K=4$, $S=10^{5}$, $\mu=10$ and $M=67$. The horizontal green line is the IS (Eq. 4) value.
Figure 8(b) shows that $r_{4}$ result is almost identical as the $r_{2}$ one, both very similar to the $r=1.0$ result. This is an strong evidence that rewiring when $\phi_{t}$ is low (lower than $\phi_{a}$) is the responsable mechanism for the observable gain on $\left\langle C\right\rangle$. Indeed, $r_{3}$ result is similar to the $r_{1}$ one, even slightly worst. The gain on the computational cost $\left\langle C\right\rangle$ can be better visualized on Fig. 9: the best result (minimum on $\left\langle C\right\rangle$) is achieved with rewiring probabilities $r=r_{2}$, $r=r_{4}$ and $r=1.0$ at $L=3$.
4.3 Discussion
The goal of the rewiring is for the target to get a neighbor with higher fitness, worth being copied. The portion of agents with $\phi>\phi_{t}$ that can become target neighbor depends on $\phi_{t}$. If the target fitness is higher, $\phi_{t}>\phi_{a}$, close to $\phi_{m}$, there are less available agents with $\phi>\phi_{t}$. So, after a rewiring, the chance of the target to get a new neighbor with higher fitness may be reduced. The rewiring in this case can be ineffective. This is the case with rewiring probabilities $r=r_{1}$ and $r=r_{3}$. On the other hand, if the target has low fitness, $\phi_{t}<\phi_{a}$, there are more agents with $\phi>\phi_{t}$ available to become target neighbor. This can make the rewiring more effective when $\phi_{t}$ is low, so that the target has a reasonable chance to get acquainted with a higher fitness agent. This is the case with probabilities $r=r_{2}$, $r=r_{4}$ and $r=1.0$.
The described reasoning can explain why probabilities $r_{1}$ and $r_{3}$ do not bring any improvement on the computational cost, regardless the landscape. Additionally, probabilities $r_{2}$ and $r_{4}$ did not bring any improvement when $K=0$. On this ease landscape, any copy is advantageous because there are no local maxima. So, when the taget (with low $\phi_{t}$) makes a copy, the rewiring does not bring additional benefit. On the difficult landscape the picture is different. Probabilities $r_{2}$ and $r_{4}$ seems to make the difference at $L=3$ as the chance do get a new model as neighbor are relevant. At lower and higher $L$, this chance seems to be reduced.
This scenario can explain why the constant rewiring $r=1.0$ has similar results as probabilities $r_{2}$ and $r_{4}$. It includes the high probability of rewiring for all values of $\phi_{t}$. If this fitness is low, close to $\phi_{s}$, the rewiring is worthwhile. If $\phi_{t}$ is high, close to $\phi_{m}$, the rewiring does not make any difference.
@@ 5 Conclusions
We chose to apply this new coevolution model on a cooperative process. Using as prototype the NK model, our results showed a convoluted behavior of the system efficiency. A natural strategy for the target is to change some of his neighbors if they have worst (smaller) fitness than him. However, if the target fitness is high, there may not be enough agents with even higher fitness to be selected with reasonable probability. So, the rewiring in this case may not be useful. In the limit, when the target is the model, it is easier if no rewiring is done, so he continue to be the model. A selfish feeling seems to reward the overal system, akin to Adam Smith invisible hand metaphor. However, if the target has a low fitness, then it is easier to find new agents with higher fitness to be selected as a new neighbor. In this case, the rewiring may be a valid strategy.
This work showed that this tunable coevolution can drastically change the behavior of the system efficiency. More studies can be interesting to study other details of this mechanism. Additionally, it opens some interesting perspectives such as the impact of this tunable coevolution on the phase diagram involving absorbing phase transition of models such as Axelrod’s, voter model and others.
@@ Acknowledgements
This work received financial support from the brazillian agencies CNPq (project 405508/2021-2) and FAPEG (project no. 401425/2023-1). This research was also supported by LaMCAD/UFG. The authors thank the anonymous referees for the valuable comments.
@@
The model was implemented using Fortran and the analysis and graphics were developed using Python (packages: Numpy [49], Matplotlib [50], NetworkX [51] and Pandas [52]) and Gnuplot. All codes and data are available upon request.
@@ A.1 NK landscape
Following the formal definition from Eq. 1, we have $f:\mathscr{D}\rightarrow\mathscr{R}$, where $\mathscr{D}$ is the set of $2^{N}$ binary vectors $\mathcal{X}$ and $\mathscr{R}$ is the real numbers set. The argument of the contributing functions $\Phi_{i}$ is given by Eq. 2. To better understand this, suppose the vector $\mathcal{Y}=(y_{0},y_{1},y_{2},y_{3})=(0,1,1,0)$ on the landscape $N=4$ and $K=2$. There will be $K+1=3$ bits on the argument of each $\Phi_{i}$. The first bit from the i-th contribution is the i-th bit of $\mathcal{Y}$, while the other bits are the $K=2$ neighbors on the right. So, the first contribution of $\mathcal{Y}$ is $\Phi_{0}(y_{0},y_{1},y_{2})=(0,1,1)$. The second contribution will be $\Phi_{1}(y_{1},y_{2},y_{3})=\Phi_{1}(1,1,0)$. On the third contribution $\Phi_{2}$, the third bit will be the fifth one of $\mathcal{Y}$. However, the string $\mathcal{Y}$ has only $N=4$ bits, so the fifth bit will be the first one. This equivalent to subtract $N$ from the position of the bit: $5-4=1$. Thus: $\Phi_{2}(y_{2},y_{3},y_{0})=\Phi_{2}(1,0,0)$. In the same way: $\Phi_{3}(y_{3},y_{0},y_{1})=\Phi_{2}(0,0,1)$ Putting together the four contributions, we have:
@@ A.2 Base 2 and 10
Conceptually, the description from previous section is enough. However, computationally, it is compelling to represent the vectors $\mathcal{X}$ and the arguments of the contributions $\Phi_{i}$ as integers. Both of them are binary sequences, 0s and 1s, numbers in base 2, and thus, can be converted to decimal numbers (base 10). In general, considering $N=4$: $(x_{0},x_{1},x_{2},x_{3})=x_{3}\cdot 2^{0}+x_{2}\cdot 2^{1}+x_{1}\cdot 2^{2}+x_{0}\cdot 2^{3}$, starting the counting from the right. Thus,